OpenStudy (anonymous):

Solve the following system by differentiating the first equation and substituting the second to get an equation in only one variable. Primes indicate differentiation with respect to t. Use the constants A, B, etc. for any arbitrary constant(s) in your solution. x'=16y y'=−9x

OpenStudy (turingtest):

I really wish you would stick around after you post so I can help you in real time.

OpenStudy (turingtest):

\[x'=16y\]\[y'=-9x\]take the derivative of the second equation wrt t and you get\[y''=-9x'\]now you can sub the expression for x' from the first equation into this last one\[y''=-9(16y)=-144y\]\[y''+144y=0\]which has the solution\[y=c_1\cos(12t)+c_2\sin(12t)\]now that we have y we can go back and solve the first equation, which is separable\[x'=\frac{dx}{dt}=16y=16c_1\cos(12t)+16c_2\sin(12t)\]\[x=16\int c_1\cos(12t)+c_2\sin(12t)dt\]\[x=\frac43[c_1\sin(12t)-c_2\cos(12t)]+c_3\]

OpenStudy (anonymous):

Sorry I had to go to work

OpenStudy (anonymous):

Also I go the same answer as you did but it's wrong :/

OpenStudy (turingtest):

@lalaly DE help please :D

OpenStudy (turingtest):

...because I am confused I see wolfram's answer and I don't get it

OpenStudy (turingtest):

it is reasonable that there should be no third constant, I was suspicious of that I tried to make a matrix of it and do the eigenvalue thing , but couldn't in short, I'm stumped

OpenStudy (anonymous):

Nevermind! i figured it out! thank you!

OpenStudy (turingtest):

how did you do it?

OpenStudy (anonymous):

x'=16y x''=16y'=16(-9x)=144x

OpenStudy (anonymous):

x''-144x=0

OpenStudy (anonymous):

r^2-144=0 r=12t

OpenStudy (anonymous):

plug that into the general equation for x

OpenStudy (anonymous):

and you get Acos(12t)+Bsin(12t)

OpenStudy (anonymous):

A and B are the same as saying c_1 and c_2

OpenStudy (turingtest):

is that all?

OpenStudy (turingtest):

but that is not the solution wolfram has and that is x, what about y ?

OpenStudy (anonymous):

for y you od the same thing with x. from the equation x'=16y you get y=x'/16 We already found that x is equal to 12 previously so no you plug it into the general equation which will giver you

OpenStudy (anonymous):

(-12/16)Asin(12t)+(12/16)Bcos(12t)

OpenStudy (anonymous):

and I check my answer it's correct

OpenStudy (anonymous):

checked*

OpenStudy (turingtest):

I did realize that answer, but it was not what wolf had also you seem to get a different answer depending on the order in which you solve the problem I still have quite a few questions, but congrads that you're done with it!

OpenStudy (anonymous):

wait I'm sorry I'm confused!

OpenStudy (anonymous):

also wolfram is not always correct it sometimes solves for something different than what you intially asked for

OpenStudy (turingtest):

ok I get it I swear I had that answer, but I doubted myself because it was so similar to the other (minus the extra constant), and different from wolf (which I know is frequently wrong about this kind of thing) oh well, thanks!